OpenStudy (anonymous):

(HELP!!) Find the values of a and b for which LMNO is a parallelogram. A.a = 50, b = 60 B.a = 60, b = 50 C.a = 50, b = 70 D.a = 70, b = 50

OpenStudy (anonymous):



OpenStudy (mayankdevnani):

HINT: opposite sides of parallelogram are equal

OpenStudy (mayankdevnani):

so, \[\huge b+20=2b-50\] solve for b???

OpenStudy (anonymous):

a+40=2a-10 -40 -40 ______________ a=2a-30 ?

OpenStudy (anonymous):

b=2b-30

OpenStudy (anonymous):

this looks like the same as your last question, with different numbers

OpenStudy (anonymous):

still confused on it. :/

OpenStudy (mayankdevnani):

\[\huge a+40=2a-10\] \[\huge -a=-50\] \[\huge a=50~degree\]

OpenStudy (anonymous):

b=60

OpenStudy (anonymous):

So the answer is A?

OpenStudy (anonymous):

just remember that opposite sides are equal in a parallelogram, set them equal to each other to find the variable

OpenStudy (mayankdevnani):

no!!!

OpenStudy (anonymous):

b=70

OpenStudy (anonymous):

it should be correct?

OpenStudy (mayankdevnani):

you are right